OpenStudy (anonymous):

In terms of n, find the sum of the first n terms of the arithmetic sequence: -4, 5, 14, 23, 32, .....

OpenStudy (anonymous):

9n-4

OpenStudy (anonymous):

srry, thats not right

OpenStudy (anonymous):

Find the difference: 32-23=9, 23-14=9, so d=9 The formula for finding the sum of an arithmetic sequence is \[S_n=\frac{n}{2}(2a+(n-1)d)\] where a is the first number (-4) and n is the number you are trying to find.

OpenStudy (anonymous):

thats no it either

OpenStudy (anonymous):

finaiized has it right

OpenStudy (anonymous):

and what would i plug in for d?

OpenStudy (anonymous):

The difference which I calculated first.

OpenStudy (anonymous):

d=9

OpenStudy (anonymous):

oh duh.

OpenStudy (anonymous):

and then just solve for n right?

OpenStudy (anonymous):

no, leave it in terms of n

OpenStudy (anonymous):

oh duh it says that in the questionnnn. I've been up too long. thanks guys!

OpenStudy (anonymous):

Yep, Sn is just the total sum, and n is the number you are trying to solve for!